. Which letter comes 2/5 of the way among A and J?
. Which letter comes 2/5 of the way among A and J?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

There are 10 letters from A to J
So, 2/5 of 10 = 2/5 × 10

=2 × 2
=4
The 4thletter from A to J is D.
Therfore, D comes 2/5 of the way among A and J.


Thanks